# Federated Continual Learning: Concepts, Challenges, and Solutions

### v1

@@ 5.1.3 Negative Knowledge Transfer
As shown in Figure 4, addressing negative knowledge transfer in FCL involves strategies such as careful design of aggregation methods, regularization techniques to preserve important knowledge, and ensuring that the learning process remains robust to the differences in data distributions across clients. These strategies are illustrated in Figure 4. To address this challenge, [63, 82] proposes FedKNOW and FedViT, respectively, which are client-side solutions for mitigating negative knowledge transfer and a novel concept of signature task knowledge to extract and integrate the knowledge of signature tasks that are highly influenced by the current task. FedKNOW is designed to continuously extract and integrate knowledge from signature tasks that significantly influence current tasks by assessing the relevance of tasks, when integrating knowledge. This assessment helps in filtering out knowledge from tasks that may not align well with the current task’s objectives. Each client in FedKNOW consists of a knowledge extractor, a gradient restorer, and, crucially, a gradient integrator. During the training of a new task, the gradient integrator plays a key role in preventing catastrophic forgetting and mitigating negative knowledge transfer. It achieves this by effectively combining the signature tasks identified from past local tasks with relevant tasks from other clients, facilitated through the global model. On the other hand, FedViT introduces a task-aware Vision Transformer architecture that is capable of distinguishing between different tasks and adapting its learning process accordingly, which reduces the risk of negative knowledge transfer by ensuring that the model’s focus remains on relevant features and patterns specific to each task, rather than indiscriminately applying knowledge from previous tasks that may not be applicable. Also, this paper utilizes the selective parameter update mechanism, which carefully chooses which parts of the model should be updated based on the new task. Moreover, The model in FedViT is designed with task-specific heads, which means that different tasks can have their own dedicated output layers. This approach reduces the likelihood of negative knowledge transfer by isolating the outputs for different tasks, ensuring that the learning from one task does not adversely affect the performance of another task. [34] addresses negative knowledge transfer by enhancing collaboration among edge devices to share relevant knowledge, employing selective knowledge sharing to filter out harmful information, and maintaining task awareness to prioritize valuable knowledge. Additionally, it uses gradient adjustment mechanisms to ensure new knowledge integrates without conflicting with prior learning, thereby enabling more robust and accurate learning across distributed edge environments.
@@ 5.2 Knowledge Dissemination
For streaming data, asynchronous updating might be favored to allow rapid adaptation, while synchronized updating might be preferred in static environments to ensure consistency across clients.
5.2.1 Updating Frequency
@@
This approach offers several advantages. Clients can send updates as soon as they are ready, reducing idle times and enhancing overall system efficiency. Additionally, asynchronous updating allows the system to scale more effectively, as it does not require synchronization across a large number of clients. The system also demonstrates greater resilience to client dropouts or delays, as it does not rely on all clients being synchronized.
@@ 5.3 Client Models Challenges
Many studies explore replay-based methods, which can be divided into direct replay and generative replay [201, 54]. These methods store a subset of past experiences or data and periodically replay them to the model, while learning new information, helping to maintain the model’s performance on previously learned tasks. Direct replay involves storing actual data samples from previous tasks and interleaving these samples with new data during the training of new tasks [54, 71, 223]. Some studies [37, 86, 66] select the most representative samples for the replay buffer. On the other hand, generative replay involves using generative models to synthesize data samples that approximate the distribution of previous tasks. Instead of storing actual data, the method generates representative samples from the learned distribution of past tasks during the training of new tasks [39, 224]. Rehearsal-based methods differ in that they generate or synthesize data representations (such as embeddings or features) to approximate the data distribution of previous tasks. Prototypical learning [87], progressive neural networks [7], and attention models [76] are used in FCL frameworks to preserve these embeddings according to their importance [64] and use them for retaining previously learned information.
Gradient-based methods aim to mitigate client forgetting by manipulating the gradient during local training in each client separately to ensure that updates made, while learning new tasks, do not interfere with the knowledge of previously learned tasks. For instance, Dong et al. [68, 46] introduce gradient compensation loss to ensure that learning new classes does not lead to forgetting old classes. Additionally, orthogonal gradient aggregation [225] ensures that updates for new tasks are orthogonal to those for old tasks by minimizing the overlap between gradient directions for different tasks.
Some papers propose optimization and loss function modifications to tackle local forgetting [5]. Split-based methods, knowledge extraction, and local optimization are used as optimization methods. For example, specific modules designed on the client side propose loss functions such as class-projection relation strengthening loss function, class-aware distribution balancing loss function [77], and proximal term loss [72] to alleviate forgetting from streaming data.
6 Furture Research Directions

### v2

@@ 5.1.3 Negative Knowledge Transfer
Table 5: Performance Comparison of FedKNOW and FedViT: Accuracy Improvement and Forgetting Rate Across Tasks and Clients on CIFAR-100
Methods Number of Clients Number of Tasks
50 Clients 100 Clients 1 2 3 4 5 6 7 8 9 10
Accuracy FR Accuracy FR Improvement (%)
FedKNOW [76] 59.26 11.30 53.18 7.30 36.52 74.74 82.58 84.84 88.69 94.87 92.40 95.52 98.72 97.75
FedViT [95] 62.08 1.70 54.25 4.23 64.6 74.22 88.3 87.5 95.0 103.03 107.69 106.81 113.74 110.66
@@ 50 Clients 100 Clients 1 2 3 4 5 6 7 8 9 10
As shown in Figure 7, addressing negative knowledge transfer in FCL involves strategies such as careful design of aggregation methods, regularization techniques to preserve important knowledge, and ensuring that the learning process remains robust to the differences in data distributions across clients. These strategies are illustrated in Figure 7. To address this challenge, [76, 95, 255] proposes FedKNOW, FedViT, and Loci respectively, which are client-side solutions for mitigating negative knowledge transfer and a novel concept of signature task knowledge to extract and integrate the knowledge of signature tasks that are highly influenced by the current task. FedKNOW is designed to continuously extract and integrate knowledge from signature tasks that significantly influence current tasks by assessing the relevance of tasks, when integrating knowledge. This assessment helps in filtering out knowledge from tasks that may not align well with the current task’s objectives. Each client in FedKNOW consists of a knowledge extractor, a gradient restorer, and, crucially, a gradient integrator. During the training of a new task, the gradient integrator plays a key role in preventing catastrophic forgetting and mitigating negative knowledge transfer. It achieves this by effectively combining the signature tasks identified from past local tasks with relevant tasks from other clients, facilitated through the global model. On the other hand, FedViT introduces a task-aware Vision Transformer architecture that is capable of distinguishing between different tasks and adapting its learning process accordingly, which reduces the risk of negative knowledge transfer by ensuring that the model’s focus remains on relevant features and patterns specific to each task, rather than indiscriminately applying knowledge from previous tasks that may not be applicable. Also, this paper utilizes the selective parameter update mechanism, which carefully chooses which parts of the model should be updated based on the new task. Moreover, The model in FedViT is designed with task-specific heads, which means that different tasks can have their own dedicated output layers. This approach reduces the likelihood of negative knowledge transfer by isolating the outputs for different tasks, ensuring that the learning from one task does not adversely affect the performance of another task. [47] addresses negative knowledge transfer by enhancing collaboration among edge devices to share relevant knowledge, employing selective knowledge sharing to filter out harmful information, and maintaining task awareness to prioritize valuable knowledge. Additionally, it uses gradient adjustment mechanisms to ensure new knowledge integrates without conflicting with prior learning, thereby enabling more robust and accurate learning across distributed edge environments.
Table 5 and Figure 8 compare the performance of FedKNOW and FedViT on the CIFAR-100 dataset across 10 tasks. The results show that FedViT consistently outperforms FedKNOW across all tasks, with notable accuracy improvements at each stage. On average, FedViT achieves a higher percentage of accuracy per task, indicating stronger generalization and better scalability to the number of tasks. Furthermore, when evaluating scalability with respect to the number of clients (50 vs. 100), both methods experience a drop in accuracy. However, FedViT maintains higher accuracy overall, while the forgetting rate (FR) decreases in FedKNOW but increases slightly in FedViT. Despite this, the overall performance of FedViT remains superior in terms of both task and client scalability.
To compare computational overhead associated with FedViT and FedKNOW across the primary phases of knowledge extraction, storage, and gradient restoration. During knowledge extraction, FedViT introduces higher computational demands by evaluating loss values across the training dataset to identify high-quality samples, whereas FedKNOW employs a more efficient weight-based pruning strategy, retaining only a small fraction (10%) of the most significant model weights. In terms of knowledge storage, FedViT maintains selected data samples, resulting in higher memory consumption, while FedKNOW achieves greater storage efficiency by preserving only the pruned model weights. Gradient restoration in FedViT requires computation over the retained sample subset, whereas FedKNOW reconstructs gradients using its stored weights in conjunction with data from the current task, reducing reliance on historical data. Both approaches employ quadratic programming to integrate gradients, incurring comparable computational complexity with polynomial time guarantees. Notably, FedViT mitigates training costs by freezing a portion of model parameters during training, while FedKNOW, despite not incorporating parameter freezing, demonstrates improved scalability and stability in training time as the retained knowledge size increases, highlighting its computational efficiency.
Figure 8: Cumulative Accuracy Improvement Comparison Between FedViT and FedKNOW on CIFAR-100: Area Under the Curve (AUC) comparison of average percentage accuracy improvement for FedViT and FedKNOW on the CIFAR-100 dataset across 10 incremental tasks. The AUC metric reflects the overall performance trend, with FedViT consistently achieving higher cumulative accuracy improvements, indicating stronger generalization and retention capabilities across tasks.
@@ 5.2 Knowledge Dissemination
For streaming data, asynchronous updating might be favored to allow rapid adaptation, while synchronized updating might be preferred in static environments to ensure consistency across clients. To better understand how different update strategies influence learning in FCL, this section examines updating frequency as a key challenge.
@@
This approach offers several advantages. Clients can send updates as soon as they are ready, reducing idle times and enhancing overall system efficiency. Additionally, asynchronous updating allows the system to scale more effectively, as it does not require synchronization across a large number of clients. The system also demonstrates greater resilience to client dropouts or delays, as it does not rely on all clients being synchronized. From a practical standpoint, [8] introduce an asynchronous federated continual learning framework in which clients learn tasks in different orders and at varying times. Their proposed method, FedSpace, tackles key challenges such as catastrophic forgetting and task misalignment, showing improved performance in asynchronous learning environments.
@@
The evaluation is conducted on CIFAR-100, partitioned into 5 class-incremental tasks to simulate realistic FCL scenarios. Two FCL settings are considered: synchronous, where all clients follow the same task sequence with varying class proportions (controlled via Dirichlet distribution, $\alpha=1$), and asynchronous, where clients have partially overlapping class distributions, some classes are globally shared while others remain private to individual clients. To examine the impact of these settings on global model performance, several incremental aggregation methods are compared, including FedAvg [3], FedProx [229], FedEWC [260], and GLFC [59] with a ResNet-18 backbone, as well as FedViT [95], FedL2P [261], FedDualP [262], and FedMGP [250] with a ViT backbone. The objective is to assess how different aggregation strategies influence global model accuracy across clients in both synchronous and asynchronous FCL settings. As shown in Figure 9, the darker bars represent the global model’s accuracy when it is updated using synchronous aggregation, while the lighter bars correspond to asynchronous updating. In the upper plot, ResNet-18 is used as the backbone model, highlighting a clear performance gap between the two aggregation strategies. Synchronous methods, such as FedAvg and FedProx, consistently achieve higher global accuracy across tasks, suggesting more stable learning and reduced forgetting. In contrast, asynchronous methods often result in lower performance, particularly in later tasks, likely due to delayed or uncoordinated client updates that disrupt the learning process. This issue is especially pronounced in methods like GLFC, where the asynchronous variant exhibits a sharp decline in accuracy, indicating significant forgetting. In contrast, the lower plot, where a Vision Transformer (ViT) is used as the backbone, demonstrates that ViT-based models maintain strong and consistent performance under both synchronous and asynchronous settings. Methods such as FedMGP and FedDualP achieve high accuracy across all tasks and exhibit minimal signs of forgetting. This robustness is attributed to ViT’s ability to capture global dependencies through self-attention mechanisms and its strong representation learning capabilities, which help preserve previously acquired knowledge even when updates are asynchronous. In summary, synchronous aggregation generally leads to more reliable performance and better retention of learned knowledge, particularly with convolutional architectures like ResNet-18. However, ViT-based models demonstrate that asynchronous aggregation can still perform competitively, making it a practical alternative in scenarios where full client synchronization is difficult to achieve.
In terms of communication efficiency, asynchronous updating is generally more efficient than synchronous methods, as it allows clients to send updates independently without waiting for others, reducing idle time and making better use of network resources. This is particularly advantageous in FCL, where data arrives incrementally and timely updates are crucial. In contrast, synchronous approaches suffer from communication bottlenecks due to client synchronization requirements, often leading to delays and underutilized bandwidth, especially in environments with heterogeneous client speeds or intermittent availability.
Figure 9: Evaluation of the aggregated global model’s accuracy on local CIFAR-100 test sets, with each client following a 5-task class-incremental learning setup.
@@ 5.3 Client Models Challenges
Many studies explore replay-based methods, which can be divided into direct replay and generative replay [213, 67]. These methods store a subset of past experiences or data and periodically replay them to the model, while learning new information, helping to maintain the model’s performance on previously learned tasks. Direct replay involves storing actual data samples from previous tasks and interleaving these samples with new data during the training of new tasks [67, 84, 263]. Some studies [50, 99, 79] select the most representative samples for the replay buffer. On the other hand, generative replay involves using generative models to synthesize data samples that approximate the distribution of previous tasks. Instead of storing actual data, the method generates representative samples from the learned distribution of past tasks during the training of new tasks [52, 264]. In the context of CL, several studies have leveraged feature representations to address the problem of catastrophic forgetting. Wang et al. [265] enhance rehearsal-based methods by incorporating loss functions based on the Hilbert-Schmidt Independence Criterion (HSIC) into the standard rehearsal framework. This approach aims to mitigate forgetting by explicitly reducing inter-task interference and promoting the learning of task-invariant features, proving particularly effective when working with limited buffer sizes. Similarly, Liu et al. [266] propose a method that combines centroid-based sampling with centroid distance distillation. This strategy reduces dataset bias during rehearsal and helps preserve essential inter-class relationships from previous tasks, thereby alleviating forgetting. Prototypical learning [100], progressive neural networks [7], and attention models [89] are used in FCL frameworks to preserve these embeddings according to their importance [77] and use them for retaining previously learned information.
Gradient-based methods aim to mitigate client forgetting by manipulating the gradient during local training in each client separately to ensure that updates made, while learning new tasks, do not interfere with the knowledge of previously learned tasks. For instance, Dong et al. [81, 59, 239, 267] introduce gradient compensation loss to ensure that learning new classes does not lead to forgetting old classes. In [81], a new method called the Local-Global Anti-forgetting model is proposed, which enhances the performance of GLFC [59]. Additionally, orthogonal gradient aggregation [268] ensures that updates for new tasks are orthogonal to those for old tasks by minimizing the overlap between gradient directions for different tasks. Furthermore, [269, 270] explore gradient-based methods specifically in CL to mitigate forgetting, these methods can be used as local settings for FCL to alleviate local forgetting. [269] proposed Knowledge Aggregation Networks (KANets), a dual-branch architecture in which one branch is frozen to preserve previously learned knowledge, while the other remains trainable to acquire new information. This structural separation mitigates gradient conflicts between new and existing knowledge, enabling a balanced trade-off between learning new data and preventing the forgetting of prior knowledge. [270] proposed Gradient Self-Adaptation (GSA) to alliviate the imbalance between the previously learnt and new data. Also, existing class incremental learning methods often assume all old classes forget uniformly, which is inaccurate and hinders effective knowledge retention. Some papers propose optimization and loss function modifications to tackle local forgetting [5]. Split-based methods, knowledge extraction, and local optimization are used as optimization methods. For example, specific modules designed on the client side propose loss functions such as class-projection relation strengthening loss function, class-aware distribution balancing loss function [90], and proximal term loss [85] to alleviate forgetting from streaming data.
5.4 Privacy Challenges of Federated Continual Learning
FCL combines the distributed, privacy-sensitive nature of FL with the dynamic, non-stationary data handling of CL. This combination makes traditional FL challenges like heterogeneity, privacy preservation, communication overhead, and resource constraints significantly more complex in the context of continuous data streams. Furthermore, FCL introduces new challenges like longitudinal data leakage and requires addressing the interaction of FL and CL issues, such as managing catastrophic forgetting at both global and local levels and mitigating negative knowledge transfer in a distributed, dynamic setting. Highlighting the need for new methodologies that effectively combine FL’s collaboration and privacy with CL’s adaptation over time is essential.
In continual scenarios, clients contribute over a period of time with evolving local data, which leads to opportunities for privacy leakage due to cumulative exposure [271, 272]. This is particularly concerning when an adversary analyzes a sequence of model updates from the same client, potentially uncovering time-sensitive patterns. For instance, even when individual updates appear harmless, their progression may reveal confidential trends such as changes in user behavior, health status, or location.
Recent research has adopted different approaches to mitigate the aforementioned risks. Temporal DP builds on traditional DP by noise injection that fluctuates over time, easing the total leakage across rounds of training [273]. Update obfuscation and compression techniques aim to conceal sensitive trends by modifying model updates in a randomized and time-dependent manner [274]. Altering client identifying markers periodically, or scheduling participations randomly aids in diminishing the association between subsequent updates made by a specific client [275]. Additionally, privacy-aware forgetting mechanisms reduce long-term exposure by down-weighting or discarding older updates [276]. These emerging solutions suggest that effective privacy preservation in FCL must consider both the distribution of data across clients and the progression of client data over time.
6 Datasets
To be able to use these techniques in contributive systems, our data requires some features to have, which are distinguished in Table 6, in which data requirements are which can be grouped into three categories. Most of the FCL research uses different publicly available datasets across various domains to validate any particular learning strategy for their challenges, including concept drift and catastrophic forgetting. Table 3.5 summarizes a comprehensive list of datasets and repositories on the basis of dimensionality, data format, type of drift, and respective repositories. The table also summarizes datasets to observe gradual and abrupt drifts with their dimensional parameters in terms of pixels, channels, or features. By providing this in-depth overview, Table 3.5 acts as a foundational point of reference for the selection of datasets that correspond to certain FCL research objectives.
7 Future Research Directions
@@
This survey examines the inherent challenges within FL, CL, and FCL, focusing specifically on those emerging from the learning process itself. Accordingly, future research should prioritize enhancing the robustness and generalization of FCL models, particularly in adversarial settings and scenarios with noise or unreliable data. Strengthening defense mechanisms against adversaries and improving model resilience to label noise and distribution shifts are essential for ensuring the reliability and security of FCL systems.
@@
Investigating advanced dynamic optimization techniques from related domains, such as those developed for Dynamic Multiobjective Optimization Problems (DMOPs) [277, 278], holds promise for addressing persistent challenges in Federated Continual Learning (FCL). In particular, concepts from Evolutionary Transfer Optimization (ETO), which aims to leverage knowledge from previous dynamic environments to accelerate adaptation in new settings, are highly relevant [277]. For example, incorporating methods like Cascaded Fuzzy Systems (CFS) to generate soft labels for historical information could enable more expressive representations of the global model’s retained knowledge [277]. Similarly, techniques such as Kernel Mean Matching (KMM), which estimate the relevance of past data by aligning data distributions without requiring explicit labels, offer a principled way to assess the applicability of prior knowledge to new client data or tasks [278]. Integrating these ideas into the FCL framework could support the development of adaptive regularization or aggregation strategies, helping to mitigate negative transfer and reduce catastrophic forgetting. Such approaches could significantly enhance the robustness and adaptability of global models operating in non-stationary, decentralized environments.